OpenStudy (anonymous):

Jackie and Parker are part of an egg drop competition at school. They each drop their egg from the top of the stairwell. Jackie drops her egg 1.5 seconds after Parker. After Jackie's egg has been falling for 4 seconds, what is the velocity of Parker's egg?

OpenStudy (shane_b):

Use the kinematic equation: \[V_{final}=V_{initial}+at\]Since they are both dropping their eggs you can assume their initial velocity is 0m/s and reduce this equation to : \[V_{final}=at\]So you end up with: \[V_{f~Parker}=(9.8m/s2)(4.0s+1.5s)=53.9m/s\]\[V_{f~Jackie}=(9.8m/s2)(4.0s)=39.2m/s\]